It is desired to have a current of 20 A in a 5- resistor when it is connected to an 80-V battery. Is there any way in which an auxiliary resistor can be connected in the circuit to increase the current in the 5- resistor to this valur? If so, what should ist resistance be?


1
Expert's answer
2021-11-03T09:50:59-0400

Assume x ohm resistor connected in parallel combination

"\frac{1}{R_{eq}}=\frac{1}{R_1}+\frac{1}{R_2}"

"R_{eq}=\frac{R_1R_2}{R_1+R_2}"

"R_{eq}=\frac{5X}{5+X}"

"V=IR_{eq}"

Put value

"80=20\times\frac{5X}{5+X}"

"20+4X=5X"

"X=20\Omega"

X=20 "\Omega" connected in parallel combination of 5"\Omega"